The monthly average salary of worker in an industry is rupees 1900. If salary of manager is also included the average salary become rupees 2000. Find the annual salary of the manager

Grade:7

3 Answers

Arun
6 years ago
Let salary of manager = x
Hence
According to question
(x + 1900)/2 = 2000
x + 1900 = 4000
x = 2100 rs.
 
Regards
Arun (askIITians forum expert)
Satyam
6 years ago
Wrong answer correct is rs 48000 thanks Arun please give me it`s notesHfbvsjemmsmmdmdkdmmsmsnznsnsnsmdkejjekejdjdjjdjdnejsjejejejdjjsjsn
Vipul kumar
6 years ago
There is a worker and a manager. so , total salary = ₹2000×2= 4000Salary of worker = 1900Salary of manager = ₹(4000 - 1900) = ₹2100